OpenStudy (anonymous):

What is the slope of the line 5x – 2y = -10?

OpenStudy (uri):

Arrange it first!

OpenStudy (anonymous):

change your equation into the form: y = mx +n then m is the slope

OpenStudy (uri):

Formula:y=mx+c

OpenStudy (anonymous):

My choices are: a.) -5/2 b.) 5/2 c.) 2/5 d.) -2/5

OpenStudy (uri):

5x – 2y = -10? -2y=-5x+10 y=-5/2x +10/2 y=-5/2x Slope.

OpenStudy (anonymous):

e.) none of the above

OpenStudy (uri):

-10*

OpenStudy (anonymous):

Okay, Thanks!